### v1

# Learning Vision-Based Neural Network Controllers with Semi-Probabilistic Safety Guarantees

Abstract
Ensuring safety in autonomous systems with vision-based control remains a critical challenge due to the high dimensionality of image inputs and the fact that the relationship between true system state and its visual manifestation is unknown. Existing methods for learning-based control in such settings typically lack formal safety guarantees. To address this challenge, we introduce a novel semi-probabilistic verification framework that integrates reachability analysis with conditional generative adversarial networks and distribution-free tail bounds to enable efficient and scalable verification of vision-based neural network controllers. Next, we develop a gradient-based training approach that employs a novel safety loss function, safety-aware data-sampling strategy to efficiently select and store critical training examples, and curriculum learning, to efficiently synthesize safe controllers in the semi-probabilistic framework. Empirical evaluations in X-Plane 11 airplane landing simulation, CARLA-simulated autonomous lane following, and F1Tenth lane following in a physical visually-rich miniature environment demonstrate the effectiveness of our method in achieving formal safety guarantees while maintaining strong nominal performance. Our code is available at https://github.com/xhOwenMa/SPVT.
I Introduction
Many real-world applications, such as self-driving cars and robotic navigation, require controllers that process high-dimensional image inputs to make real-time decisions. The centrality of visual inputs (particularly when other modalities are limited or unreliable) thus makes ensuring the safety of vision-based control an important problem in trustworthy AI. However, verifying the safety of such controllers remains a major open challenge due to the complexity of image-based inputs and the high computational cost of traditional verification methods [1, 2, 3].
While reinforcement learning (RL) with high-dimensional image inputs has shown promise in learning control policies that optimize performance [4, 5], most methods lack formal guarantees of safety [6]. Moreover, verifying the safety of neural network controllers operating in high-dimensional observation spaces remains computationally intractable. Existing approaches to safe control primarily focus on low-dimensional state inputs [7] and empirical safety evaluations. Other approaches, such as verified safety over the entire input region and control barrier function-based methods, have also been explored [8, 9]. However, these methods struggle when the controller operates on image inputs due to the high dimensionality of the observation space. Moreover, while dynamic behavior of many autonomous systems of interest has established models approximating their trajectories through the system state space, physics that map state to its visual representation are considerably more complex, and associated models far more involved and less reliable.
In this work, we integrate reachability analysis with generative modeling to enable efficient verification of neural network controllers operating on high-dimensional image spaces [10, 11]. Specifically, we employ a conditional generative adversarial network (cGAN) [12] to model the perceptual mapping from states to images, allowing us to verify safety properties in a structured and lower-dimensional latent space. To address the scalability challenges of verification, we introduce a semi-probabilistic verification (SPV) framework, where safety properties are verified over a sampled distribution of initial states (using distribution-free tail bounds) but for all possible latent environment representations of the cGAN. In addition, we present a training algorithm that makes use of a novel safety loss as a differentiable proxy to this verification objective. A key component of this algorithm is our approach to adapt the training set, stochastically biasing it towards states for which safety is difficult to verify. As our experiments in simulated plane landing, as well as both simulated and physical autonomous lane following, demonstrate, the proposed approach yields control policies that exhibit considerably stronger safety properties compared to state-of-the-art safe vision-based control baselines.
In summary, our key contributions are as follows:
  • A novel semi-probabilistic safety verification framework (SPV) approach that provides formal safety guarantees while remaining computationally feasible in high-dimensional vision-based control settings.
  • A novel training approach which uses a differentiable proxy loss for SPV and maintains a dynamic training set which adaptively prioritizes safety-critical states.
  • Experimental evaluation in simulated and physical lane following that demonstrates effective empirical and provably-verified performance of the policies trained through our approach in comparison with several state-of-the-art baselines.
II Related Work
Formally verifying the safety of vision-based controllers is very challenging. Traditional verification tools [2, 3, 13, 14] are too computationally demanding to scale to realistic networks. Recent progress have enabled verification of larger neural networks through over-approximation [15, 16, 17], and abstract interpretation [18, 19, 20]. However, verifying vision-based controllers remains a difficult task since it is inherently less straightforward to define safety properties in the image space.
A practical approach to verify vision-based controllers is by approximating the perception module. This reduces verification complexities by projecting the problem from image space to lower-dimensional state space [10, 11, 21]. These methods, however, only focus on post-hoc verification and cannot be integrated into the training process for learning a safer controller. For a more comprehensive survey on the verification of vision-based controllers, see [22].
Safe reinforcement learning encodes safety based on the constrained Markov decision processes [23] but it is often too soft to enforce strict safety constraints [24]. On the other hand, control-theoretic methods such as barrier functions [7, 9, 25], and reachability analysis [26, 27] are more powerful but they have limited scalability to high-dimensional system such as vision-based controllers. Recently, some works extend these methods to the image space [28, 29, 30, 31]. However, they function more like a safety check around the controller, giving yes/no answers for passed in reference control, without providing formal safety guarantees about the controller itself.
III Model
III-A Problem Formulation
We consider a discrete-time dynamical system:
$s_{t+1}=f(s_{t},u_{t}),\ \ o_{t}=h(s_{t},\omega),\ \ s_{0}\sim\mathcal{D},\ \ \omega\sim\Omega,$ (1)
where $s_{t}\in\mathcal{S}$ is the system state (e.g., position, steering angle of the vehicle), $o_{t}\in\mathcal{O}$ is the vision-based (image) observation perceived by the agent, $u_{t}\in\mathcal{U}$ the control action, $h$ the mapping from state to observation, and $\mathcal{D}$ a distribution over the initial state $s_{0}$. Notably, $h$ takes as input a perceptual environment $\omega$, which models an unobserved source of environment variation distributed according to an unknown distribution $\Omega$. We assume that the dynamics $f$ are known (for example, well-known dynamical system models for physical systems), while $h$ and $\mathcal{D}$ are both unknown. At execution time, we suppose that only observations $o_{t}$ are known to the controller, with state $s_{t}$ unobservable. Our goal is to synthesize a control policy $\pi$ mapping visual observations $o$ to control actions $u$ which is provably safe in the sense we formalize next.
Let $P$ denote a safety specification, which is a predicate $P(s)$ indicating whether a state $s\in\mathcal{S}$ is safe or not. Similarly, let $P(\mathcal{T})$ indicate whether $P(s)$ is true for all $s\in\mathcal{T}\subseteq\mathcal{S}$. We assume that both $P(s)$ and $P(\mathcal{T})$ can be evaluated efficiently (for example, safety is often described using linear inequalities and $\mathcal{T}$ is a polyhedron). Given a policy $\pi$, the controlled dynamical system effectively becomes $s_{t+1}=f(s_{t},\pi(h(s_{t}))$. We say that this dynamical system is safe for an initial state $s_{0}$ over a horizon $K$ if $P(s_{t})$ is true for all $0\leq t\leq K$.
This notion of safety, however, is limited for two reasons. First, we do not know $h$, so we cannot directly verify the dynamical system above. Second, we wish for a policy $\pi$ to satisfy safety in a way that is not tied to a specific starting state, but with respect to the full set of initial states $\mathcal{S}$. We address the first challenge by leveraging a conditional generative adversarial network (cGAN) to approximate $h$, and the second by using a semi-probabilistic verification (SPV) framework. We describe both of these ideas next.
III-B Approximating the Visual Observation Model
We address the first challenge by using a conditional generative model $g(s,z)$ which induces a distribution over observations $o\in\mathcal{O}$ for a given state $s\in\mathcal{S}$, with $z\in\mathcal{Z}$ a (typically uniformly distributed) random vector, analogous to the approach proposed by Katz et al. [10]. Such a generator can be trained, for example, using the conditional generative adversarial network (cGAN) framework [12, 32] from a collection of data $(o,s)$ in which images $o$ are annotated with associated states $s$. We can view the latent random vectors $z$ as representations of natural environment variation (e.g., different perspectives, lighting, etc). The goal here is that $g$ approximates $h$, but in practice this assumption is too strong. Instead, we make the following considerably weaker assumption about the relationship between $h$ and $g$.
Assumption 1
$\sup_{s,\omega}\inf_{z}\|h(s,\omega)-g(s,z)\|\leq\epsilon$.
In practice, this assumption boils down to having (a) sufficient training data for the generator $g$ and (b) a sufficiently rich representation (e.g., neural network) and latent dimension of $z$.
IV Semi-Probabilistic Verification
Our notion of safety is based on $K$-reachability. The principal distinction is that in the latter, safety is guaranteed for all states in some specified set $\mathcal{S}_{0}\subseteq\mathcal{S}$ from which dynamics may be initialized. However, the resulting $K$-reachability proofs are generally conservative and typically suffer from significant scalability challenges. When controllers use vision, scalability can be a prohibitive barrier to verification. In practice, however, we can often obtain information about the distribution over initial states of the dynamical system $s_{0}$. For example, by collecting empirical visual data and annotating it with state-relevant information (for example, Waymo [33] or KITTI [34] datasets in the case of autonomous driving). On the other hand, the distribution over the initial state is often difficult to cleanly characterize (indeed, it may be heavy-tailed). It is, therefore, natural to appeal to distribution-free bounds to obtain probabilistic safety proofs with respect to the unknown distribution $\mathcal{D}$ over $s_{0}$ based on safety properties obtained for a finite sample of initial states. In contrast, the distribution of the visual environment induced by $\omega$ is far more challenging to characterize or sample, particularly since we do not know $h$.
We propose to balance these considerations through a semi-probabilistic verification (SPV) framework, in which we aim to obtain provable distribution-free guarantees with respect to $\mathcal{D}$, but which hold in the worst case with respect to environment variation $\omega$.
To formalize, fix a policy $\pi$ and let $\mathcal{S}_{t+1}(s_{0},\pi)=\{f(s,\pi(o))|o=h(s,\omega),s\in\mathcal{S}_{t}(s_{0},\pi),\omega\in\Omega\},$ where $\mathcal{S}_{0}(s_{0},\pi)=\{s_{0}\}$. Define $\mathrm{Reach}_{K}(s_{0},\pi)=\cup_{t=0}^{k}\mathcal{S}_{t}(s_{0},\pi)$, that is, all states that can be reached from $s_{0}$ for any perceptual environment $\omega\in\Omega$. Note that this form of reachability cannot be verified, since we do not know $h$. However, we can now leverage the cGAN $g$ as a proxy, with Assumption 1 allowing us to obtain sound safety guarantees. Specifically, let
$\hat{\mathcal{S}}_{t+1}(s_{0},\pi)=\{f(s,\pi(o))|$ $o\in g(s,z)\pm\epsilon,$
$s\in\hat{\mathcal{S}}_{t}(s_{0},\pi),z\in\mathcal{Z}\},$
and define $\mathrm{Reach}_{K}(s_{0},\pi,g)=\cup_{t=0}^{k}\hat{\mathcal{S}}_{t}(s_{0},\pi)$. The following result allows us to focus on verification with respect to $\mathrm{Reach}_{K}(s_{0},\pi,g)$.
Theorem IV.1
Under Assumption (1), $\mathrm{Reach}_{K}(s_{0},\pi)\subseteq\mathrm{Reach}_{K}(s_{0},\pi,g)$. Therefore, $P(\mathrm{Reach}_{K}(s_{0},\pi,g))\Rightarrow P(\mathrm{Reach}_{K}(s_{0},\pi))$.
Proof:
Suppose $\mathcal{S}_{t}\subseteq\hat{\mathcal{S}}_{t}$ and let $s\in\mathcal{S}_{t}$ and $o=h(s,\omega)$ for some $\omega\in\Omega$. Then $s\in\hat{\mathcal{S}}_{t}(s_{0},\pi)$ and by Assumption (1), $o\in g(s,z)$ for some $z\in\mathcal{Z}$. Consequently, $f(s,\pi(o))\in\hat{\mathcal{S}}_{t+1}(s_{0},\pi)$. Since $\hat{\mathcal{S}}_{0}(s_{0},\pi)=\mathcal{S}_{0}(s_{0},\pi)=\{s_{0}\}$, the result follows by induction. ∎
In practice, we will make use of a verification tool that is able to efficiently obtain an over-approximation of $\mathrm{Reach}_{K}(s_{0},\pi,g)$, which maintains soundness.
Our next step is to combine this with a distribution-free tail bound with respect to the initial state distribution $\mathcal{D}$. Specifically, suppose that we have a finite sample of $N$ initial states $\{s_{i}\}_{i=1}^{N}$ i.i.d. from $\mathcal{D}$. Next we show that by verifying only with respect to this finite sample of $N$ states, we can achieve a semi-probabilistic safety guarantee for the entire initial region with respect to the unknown distribution $\mathcal{D}$.
Theorem IV.2
Suppose that $\{s_{i}\}_{i=1}^{N}$ i.i.d. from $\mathcal{D}$ and let $V=\{s_{i}|P(\mathrm{Reach}_{K}(s_{i},\pi,g))\}$. Then under Assumption (1),
$\Pr_{s\sim\mathcal{D}}\left[P(\mathrm{Reach}_{K}(s,\pi))\right]\geq\frac{|V|}{N}-\sqrt{\frac{1}{2N}\log\frac{2}{\delta}}$
with probability at least $1-\delta$
Proof:
Let $\alpha=\Pr_{s\sim\mathcal{D}}\left[P(\mathrm{Reach}_{K}(s,\pi,g))\right]$ and $\hat{\alpha}=\frac{|V|}{N}$. By the Chernoff-Hoeffding bound, $\Pr\left(|\hat{\alpha}-\alpha|\geq\epsilon\right)\leq 2e^{-2N\epsilon^{2}}$, where the probability is with respect to datasets of $N$ initial states. Letting $\delta=2e^{-2N\epsilon^{2}}$, we obtain the confidence bound: $\Pr\left(\alpha\geq\hat{\alpha}-\sqrt{\frac{1}{2N}\log\frac{2}{\delta}}\right)\geq 1-\delta$. Finally, since by Theorem IV.1, $P(\mathrm{Reach}_{K}(s_{0},\pi,g))\Rightarrow P(\mathrm{Reach}_{K}(s_{0},\pi))$, the result follows. ∎
The SPV framework above can thereby combine reachability over a finite sample of initial states to yield a rigorous tail bound guarantee for safety over a given safety horizon $K$. This, of course, is for a given policy $\pi$. In the next section, we turn to the main subject of our work: synthesizing control policies $\pi$ for the dynamical system (1) that achieve strong semi-probabilistic guarantees of this kind.
V Learning-Based Synthesis of Provably Safe Vision-Based Control
At the high level, our goal is to learn a policy $\pi$ which has a long safety horizon $K$ (that is, does not reach an unsafe state for any possible trajectory over as long a horizon $K$ as possible) with high probability $1-\delta$. Suppose that $\pi_{\theta}$ is parametric with parameters $\theta$ (e.g., a neural network), and $K$ (i.e., the target safety horizon) is fixed. Our goal is to maximize the probability that a trajectory is safe for at least $K$ steps, that is,
$\max_{\theta}\Pr_{s\sim\mathcal{D}}\left[P(\mathrm{Reach}_{K}(s,\pi_{\theta}))\right].$ (2)
To make this practical, we can only rely on a finite sample of initial states, as well as make use of the cGAN $g$. Consequently, the revised proxy objective is
$\max_{\theta}\sum_{i}P(\mathrm{Reach}_{K}(s_{i},\pi_{\theta},g)).$ (3)
The previous section shows that this still enables rigorous semi-probabilistic verification. Additionally, we consider a special case in which safety properties are tied to a scalar safety score (for example, cross-track error in lane following). In particular, let $\sigma(\mathrm{Reach}_{K}(s_{i},\pi_{\theta},g))$ be a safety score function over the reachable set, with $P(\mathcal{T})$ translating the safety score into a predicate (e.g., error exceeds a predefined threshold). We assume that we can obtain differentiable bounds on $\sigma(\mathrm{Reach}_{K}(s_{i},\pi_{\theta},g))$ (e.g., if we use $\alpha,\beta$-CROWN [15, 17, 35]).
V-A The Learning Framework
A central challenge in synthesizing a provably safe policy $\pi_{\theta}$ in our setting arises from the involvement of high-dimensional images generated by the generator $g$ (as a proxy for the true perception model $h$), which serve as inputs to the controller. Our overall approach is as follows. First, we begin with a controller $\pi_{\hat{\theta}}$ that is empirically safe (e.g., pre-trained with a safe RL method), which we also use as an anchor controller to avoid sacrificing too much empirical performance as we train for safety verification. Next, starting with $\hat{\theta}$, we train (or fine-tune) $\pi_{\theta}$ to minimize $\sum_{i\in S_{l}}\mathcal{L}(s_{i},\theta),$ where $S_{l}$ is a set of initial states $s_{0}$ used in training which evolves over training iterations $l$ and $\mathcal{L}(\theta)$ an appropriate loss function. The central algorithm design questions thus amount to 1) the choice of the loss function, and 2) the problem of selecting data $S_{l}$ to use for training in each iteration, so as to ultimately obtain a provably (rather than merely empirically) safe policy. We address these questions next.
V-B Loss Function
We propose a loss function that integrates both a supervised learning loss and safety loss as follows:
$\mathcal{L}(s_{i},\theta)=\lambda_{1}\mathcal{L}_{acc}(s_{i},\theta)+\lambda_{2}\mathcal{L}_{safety}(s_{i},\theta).$ (4)
The supervised loss $\mathcal{L}_{acc}$ is the mean squared error loss between predictions of our controller and the anchor controller, and aims to preserve a strong empirical performance with respect to the original pre-trained anchor controller.
Turning next to the safety loss, recall that we assume that safety is quantified by a safety score function $\sigma(\mathrm{Reach}_{K}(s_{i},\pi_{\theta},g))$. One candidate would simply be to use this score as part of the loss function. However, this is impractical, as it is typically intractable to compute at scale and to the extent that it can be done, the tools for doing so are not differentiable. However, neural network verification techniques exist which compute differentiable sound upper and lower bounds on this quantity, and these therefore make natural candidates to use in constructing a loss function. More precisely, let $\underline{\sigma}(\mathrm{Reach}_{K}(s_{i},\pi_{\theta},g))\leq\sigma(\mathrm{Reach}_{K}(s_{i},\pi_{\theta},g))\leq\overline{\sigma}(\mathrm{Reach}_{K}(s_{i},\pi_{\theta},g))$ (i.e., $\underline{\sigma}(\cdot)$ is the lower and $\overline{\sigma}(\cdot)$ the upper bound on $\sigma(\cdot)$). To simplify notation, we let $\sigma^{(i)}_{K}=\sigma(\mathrm{Reach}_{K}(s_{i},\pi_{\theta},g))$, with $\overline{\sigma}^{(i)}_{K}$ and $\underline{\sigma}^{(i)}_{K}$ the corresponding upper and lower bounds. Then we can define the safety loss as
$\mathcal{L}_{safety}(s_{i},\theta)=\frac{|\overline{\sigma}_{K}^{(i)}|+|\underline{\sigma}_{K}^{(i)}|}{K-1},$ (5)
which measures the rate of change of the reachable region.
V-C Adaptive Training Data
Our adaptive training procedure performs gradient updates by sampling batches from an adaptive collection of training data $S_{l}$ which consists of two disjoint and fixed-size components: the set of random initial states $S_{0}$, and $S_{A}$, maintained as a priority queue, containing initial states for which safety is a challenge to satisfy. Specifically, when training starts, $S_{A}$ is empty, and we gradually populate $S_{A}$ during the warmup period by adding the $m$most challenging datapoints (in the sense detailed below) from each training batch to it. To ensure $S_{0}$ and $S_{A}$ remain disjoint, whenever a datapoint is added to $S_{A}$, it is also deleted from $S_{0}$, and we generate another datapoint uniformly randomly to add to $S_{0}$.
We select datapoints to add to $S_{A}$ based on the rate of change in the safety margin $\sigma_{i}$ over an entire $K$-step trajectory. For example, a datapoint where the vehicle deviates from the center of the lane and drifts toward the margin at high speed is prioritized. This allows us to detect points that are likely to become unsafe ahead of time. The safety loss defined in Equation (5) can be directly used as this metric here. If training fails to improve safety on datapoints in $S_{A}$, $S_{A}$ effectively becomes fixed once full. Otherwise, if safety improves or we encounter new datapoints that are less safe, $S_{A}$ adapts to reflect such changes.
After we have enough datapoints in $S_{A}$, future training batch $T$ with size $L$ consists of $\lfloor p\cdot L\rfloor$ datapoints from $S_{A}$ and the rest from $S_{0}$, where $p$ is a tunable parameter during training. The portion of datapoints from $S_{0}$ can be sampled uniformly. The portion of datapoints from $S_{A}$ are sampled following the Efraimidis & Spirakis [36] weighted sampling approach to prioritize datapoints that are more difficult (less safe) than others. Specifically, for each datapoint $i$ in $S_{A}$, we first calculate a safety parameter
$z^{(i)}$ $=\mathcal{L}_{safety}^{(i)}$ (6)
$+\gamma\Big{(}\max(0,|\overline{\sigma}^{(i)}_{K}|-\beta)+\max(0,|\underline{\sigma}^{(i)}_{K}|-\beta)\Big{)},$
where the second term further penalizes datapoints whose upper and lower bound is outside of the $[-\beta,\beta]$ region, where $\beta$ is a predefined safety threshold, and ensure such difficult-to-verify inputs are more likely to be sampled. We then assign a weight $w_{i}=e^{\alpha\cdot z^{(i)}}$ (where $\alpha$ is a hyperparameter) to each datapoint $i$ and normalize the weights such that $\sum_{i=1}^{N}w_{i}=1$. Finally we assign each element $x_{i}$ a key $k_{i}=w_{i}^{1/U_{i}}$, where $U_{i}\sim\text{Uniform}(0,1)$. Selecting the top $L$ elements with the highest $k_{i}$ values yields a weighted random sample without replacement, where each element $x_{i}$ is included with probability: $P(x_{i}\in S)=w_{i}/\sum_{j\in\mathcal{B}}w_{j}$.
The full algorithm for data sampling is shown in Algorithm 1. Line $3$ calculates how many datapoints to sample from the buffer based on a tunable parameter $p$; line $4-8$ computes the weights and keys following [36]; line $9-11$ construct the training batch. This approach enables us to have an adaptive training set which maintains the controller’s average performances while improving its verifiability with respect to datapoints that are more difficult.
Algorithm 1 Data Sampling Algorithm
1:  Input: $S_{l}=\mathcal{S}_{0}\cup\mathcal{S}_{A}$, $\beta$, $\alpha$, $p$, $L$
2:  Output: Sampled points $T$
3:  $L_{A}\leftarrow\lfloor p\cdot L\rfloor$; $L_{0}\leftarrow L-L_{A}$
4:  for each datapoint $x_{i}\in\mathcal{S}_{A}$ do
5:     Compute $z^{(i)}$ according to Equation 6
6:     Compute weight $w_{i}=e^{\alpha\cdot z^{(i)}}$
7:  end for
8:  Normalize $w_{i}$ and compute keys $k_{i}$ for all $x_{i}\in S_{A}$
9:  $T_{A}\leftarrow$ top $L_{A}$ elements in $S_{A}$ by $k_{i}$ values
10:  $T_{0}\xleftarrow{}$ uniformly sample $L_{0}$ points from $\mathcal{S}_{0}$
11:  $T\leftarrow T_{A}\cup T_{0}$
12:  Return $T$
V-D Curriculum Learning
Finally, as in prior work [8], we use curriculum learning. Specifically, during training, we first target $K_{i}$-step verified safety. After a series of epochs, we progress to $K_{i+1}$-step verified safety, where $1\leq K_{1}<K_{2}<\cdots<K_{n-1}<K_{n}=K$. This gradual increase in the verification horizon helps improve training stability and enables the controller to learn more complex safety constraints.
VI Experiments
VI-A Experiment Setup
We evaluate our approach in three settings: 1) the X-Plane 11 Flight Simulator, 2) the CARLA Simulator, and 3) a mini-city miniature urban physical autonomous driving platform with an F1Tenth racing car. We use two evaluation metrics: 1) empirical performance and 2) lower bound probability for safety guarantee as a function of $K$. We use three baselines for comparison: 1) RESPO [27], a safe reinforcement learning framework using iterative reachability analysis; 2) SAC-RCBF [9], which incorporates safety as a robust-control-barrier-function layer into training; and 3) VSRL [8], which guarantees finite-horizon safety by integrating incremental reachability verification into safe reinforcement learning.
Specifically, we first consider the autonomous aircraft taxiing problem using the X-Plane 11 Flight Simulator [37]. For training the image generator, we collected 20,000 state-image pairs. Each sample consists of an RGB image captured by a forward-facing camera, the corresponding state information (lateral offset $d$ ranging from $-10$ to $10$ meters, heading error $\theta$ between $-0.5$ and $0.5$ radians), and control inputs (nose wheel steering angle $\gamma$). Data was collected while the aircraft operated at a constant ground speed.
Our second set of experiments consider autonomous lane following using the CARLA Simulator [38] version 0.9.14. For training the image generator, we sample initial states consisting of the lateral distance $d$ from the lane center, the heading error $\theta$ relative to the lane direction, and the global coordinates $(x,y,z)$ within the CARLA map. The dataset includes trajectories with $d\in[-0.8,0.8]$ meters and $\theta\in[-0.15,0.15]$ radians. Final dataset contains 20,000 state-image pairs collected across different towns (maps) and environmental conditions.
Finally, we evaluated our approach using the F1Tenth racing car, an open-source 1:10 scale autonomous vehicle, for lane following in the mini-city physical testbed. The F1Tenth vehicle is equipped with a front-facing camera that provides visual input for lane following and can achieve scaled speeds comparable to full-scale autonomous vehicles. This platform enables us to evaluate the transferability of our training framework to real-world physical systems. We collected $400$ images in the mini-city and manually annotated the state information for these. We then finetuned the image generator from the CARLA experiments on this dataset to obtain the image generator for the F1Tenth experiments.
For all experiments, we pretrain an anchor controller, which is a dense neural network mapping images to continuous control output. The training is done by imitation learning from a tuned PID controller for the corresponding task, with learning rate of $0.0005$, batch size $256$, and $200$ epochs.
VI-B Image Generator Training and Evaluation
The image generators are implemented as conditional GAN that maps from state information and a dimension $10$ latent vector to greyscale images. The latent vector $z$ is sampled from a uniform distribution $\mathcal{U}(-1,1)$, aiming to capture semantic variations in the driving scene that are not explicitly represented in our state information (e.g., lighting conditions, road textures, environmental elements). The generators output $8\times 16$ greyscale image.
We train the GAN in two stages. First, we train a convolutional GAN. We apply spectral normalization [39] to all discriminator layers and orthogonal regularization [40] to the generator loss function. Both generator and discriminator networks are initialized using orthogonal initialization. They are trained with batch size $128$, learning rate 7e-4, and for $100$ epochs. Then, we distill this GAN into a smaller MLP generator. This simpler network architecture reduces the verification overhead significantly while preserving the state-to-image mapping. During distillation, we minimize a combination of losses $\mathcal{L}_{L1}+\lambda_{adv}\mathcal{L}_{adv}$ where $\mathcal{L}_{L1}$ is the $\ell_{1}$ distance between images generated by the teacher and student generators, $\mathcal{L}_{adv}$ is the adversarial loss from the pre-trained convolutional discriminator.
VI-C Results
Empirical Performance: To evaluate the controllers’ empirical performance, we simulate trajectories starting from $100$ random initial states for $200$ steps. We use the undiscounted cumulative rewards $\sum r_{t}$ as the evaluation metric, where the reward at each step $t$ is defined as $r_{t}=1-\min(1,|d_{t}|/\beta)$, with $d_{t}$ the cross-track error at step $t$ and $\beta$ being a predefined safety threshold.
Figure 1: Empirical performance comparison of controllers.
As shown in Figure 1, the proposed SPVT approach maintains the controller’s empirical performances in all three experiment settings, and it is generally on par with RL controllers in terms of reward maximization.
Safety Guarantee: We collected a dataset of $2000$ i.i.d. initial states. For verification, we used $\alpha,\beta$-CROWN [17, 15]. The safety property we consider is the vehicle not leaving the current lane, equivalently, $|d_{i}|\leq\beta$ for $0\leq i\leq K$. After verifying the controller on this dataset, we can calculate the lower confidence bound for the entire initial state region following Theorem IV.2.
Figure 2: Semi-Probabilistic Verification (SPV) results: $x$-axis marks the target verification trajectory length ($K$). $y$-axis is the lower bound probability of safety guarantees.
Figure 2 shows the tail bound for safety guarantee over a given range of safety horizon $K$. Compared to all the baseline methods, the proposed SPVT approach significantly increases this probabilistic bound.
VII Conclusion and Limitations
We introduced a semi-probabilistic verification framework for efficiently training and verifying vision-based neural network controllers. Our method models the perceptual mapping from state to image with a conditional GAN and uses distribution-free tail bounds to get safety guarantees over the entire initial state space. We designed a differentiable proxy to the safety verification objective under the SPV framework that can be directly incorporated into gradient-based training, and an adaptive training set that prioritizes states for which safety property is difficult to verify. While our experiments demonstrate the efficacy of our approach, many limitations remain. For the moment, our controllers and image generator require images to be grayscale and relatively low resolution, and latent dimension of the generator is relatively small. Further research is needed to handle high-resolution visual inputs, as well as to extend to multi-modal sensing.